Changing k_{f}

April 23, 2022, Revision: 1.8

1

Hypothesis:

Ω_{GW} = (qΩ_{M}/k_{f})^{2}  (1)

Determine q; we find q ∝ k_{f}^{1/4}; see Figure 2. The GW energy is high by comparison with earlier work; see Figure 3.

Figure 1: Black, blue, and red are for k_{f} = 6, 20, and 60, respectively.

Table 1: All runs with η = 2 × 10^{−3}.

k_{1}  f_{0}  Ω_{M}  Ω_{GW}
1 6.0e-02 3.874e-02 4.189e-05
3 3.2e-02 4.038e-02 2.313e-05
10 2.3e-02 3.831e-02 1.375e-05

Figure 2: Ω_{GW} = 5.2 × 10^{−4} (Ω_{M}/k_{f})^{1/2}

Figure 3: Positions of the new runs in a diagram showing Ω_{GW}^{sat} versus Ω_{M}^{max}, For orientation the old data points of the Ref. [1] are shown as gray symbols.

References

[1] A. Roper Pol, S. Mandal, A. Brandenburg, T. Kahniashvili and A. Kosowsky, “Numerical Simulations of Gravitational Waves from Early-Universe Turbulence,” Phys. Rev. D 102, 083512 (2020). doi:10.1103/PhysRevD.102.083512

                                                                                         $Header: /var/cvs/brandenb/tex/emma/k1GW/notes.tex,v 1.8 2021/01/24 05:18:52 brandenb Exp $